OpenStudy (babyslapmafro):

Two particles, A and B, are in motion in the xy-plane. Their coordinates at each instant of time t(t is greater than or equal to 0) are given by the following (click to see). Find the minimum distance between A and B...

OpenStudy (babyslapmafro):

\[x _{A}=t, y_A=4t, x_B=13-t, y_B=t\]

OpenStudy (anonymous):

Difference of x = |13-2t| Difference of y = |3t| Pythagorean Theorem \[Distance = \sqrt{(13-2t)^2+(3t)^2}= \sqrt{13t^2-52t+169}=\sqrt{13(t-2)^2+117}\] Since (t-2)^2 is always bigger or equal to 0, Distance is bigger or equal to \[\sqrt{117}\] Therefore, minimum distance is \[\sqrt{117}=3\sqrt{13}\]